OpenStudy (anonymous):

What is an equation for the translation of x^2+y^2=10 by 3 units right and 2 units down? Please show all steps!

OpenStudy (anonymous):

@Mertsj

OpenStudy (mertsj):

Do you recognize what that equation would be if you graphed it?

OpenStudy (anonymous):

A smaller circle?

OpenStudy (mertsj):

It would be a circle. What is its center?

OpenStudy (anonymous):

0,0 I believe

OpenStudy (mertsj):

Very good!!! So if you moved that center 3 units to the right and 2 units down, what would its new center be?

OpenStudy (anonymous):

(3,-2)... I think?

OpenStudy (mertsj):

Yes. You're on a roll. Can you write that same equation with the new center?

OpenStudy (anonymous):

Ooh yay! I had no idea I could do this lol. And hang on, let me try.

OpenStudy (anonymous):

3^2+(-2)^2=10

OpenStudy (mertsj):

Close. what happened to the x and y?

OpenStudy (mertsj):

\[(x-3)^2+(y+2)^2=10\]

OpenStudy (mertsj):

You have to have the x and y because they represent all the gazillions of points that make up the circle.

OpenStudy (anonymous):

Oh okay, that makes sense. But why are the signs switched in the new equation? If the points are 3 and -2, shouldn't the equation be x+3 and y-2 instead?

OpenStudy (mertsj):

Because the general equation of a circle with center (h,k) and radius r is:

OpenStudy (mertsj):

\[(x-h)^2+(y-k)^2=r^2\]

OpenStudy (mertsj):

So if you put 3 where the h is and -2 where the k is you will get the equation I posted above.

OpenStudy (anonymous):

Oh ookay. The second part got switched to a positive because it was already a negative

OpenStudy (mertsj):

Bingo!!

OpenStudy (anonymous):

Thanks so much for your help!!! :)

OpenStudy (mertsj):

yw